OpenStudy (anonymous):

Simplify: (5x^2 + 3x + 4) + (2x^2 − 6x + 3) 7x2 + 9x + 7 7x2 + 3x + 7 7x2 − 3x + 1 7x2 − 3x + 7

OpenStudy (anonymous):

Factorize?

jimthompson5910 (jim_thompson5910):

5x^2 + 2x^2 = ???

OpenStudy (anonymous):

\[ \left(2 x^2-6 x+3\right)+\left(5 x^2+3 x+4\right)=7 x^2-3 x+7 \]

OpenStudy (anonymous):

Just add them up

OpenStudy (anonymous):

I imagine that your goal now is make it into binomial?

OpenStudy (anonymous):

like \[(a+b)^2\]

OpenStudy (anonymous):

Just add them up. Thant is all what you have to do.

OpenStudy (anonymous):

thanks !